A landscaping company is hired to mow the grass for several large properties. The total area of the properties combined is 1,350 acres. The rate at which one person can mow is as follows: ŷ = 1350 − 1.5x where x is the number of hours and ŷ represents the number of acres left to mow. How many hours will it take to mow all of the lawns? (When is ŷ = 0? Enter an exact number as an integer, fraction, or decimal.)
Answer #1

ŷ = 1350 - 1.5%. - given 9 = 0. 1350- 1:5 x = 0, 1.5X = x = 1350 1350 1.5 x= goo hours it will mow all take to of the lawn